Chapter 4 CostVolumeProfit Relationships EXERCISE -13 Missing Data Base Cost-Volume-Profit Concepts ILOILO Fill in the missin Question #4-13


Salement LOLLO Castro armato BU ROSE Total Permit Son Fisad e s... Required Prepare a new contribution format income statemen Question #4-12
Answer #1

According to the requirement, first we have to solve Question #4-13

Requirement a-: Assume that only one product is being sold in each of the following four Situations:

Case Units Sold Sales Variable Expenses Contribution Margin Per Unit Fixed Expenses Operating Income (Loss)
1. 15,000 $180,000 $120,000 $4 (Working Note) $50,000 $10,000 (Working Note)
2. 4,000 (Working Note) $100,000 $60,000 (Working Note) $10 $32,000 $8,000
3. 10,000 $200,000 (Working Note) $70,000 $13 $118,000 (Working Note) $12,000
4. 6,000 $300,000 $210,000 (Working Note) $15 (Working Note) $100,000 $(10,000)

Working Note:-

Case 1. Calculation of Contribution Margin per unit.

Solution:- There is a formula to calculate contribution margin per unit i.e. Sales per unit - Variable cost per unit

Contribution margin per unit = $12 -$8 = $4 per unit

Sales = $180,000 / 15000 units = $12 per unit

Variable expenses = $120,000 / 15000 units = $8 per unit

# Calculation of Operating Income (loss).

Solution:- Sales - Variable Expenses = Contribution Margin - Fixed Expenses = Operating Income (Loss)

= $180,000 - $120,000 = $60,000 -$50,000 = $10,000 Operating Income

Case 2. Calculation of Variable Expenses.

Solution :- First of all, we have to find the Contribution Margin

Contribution Margin - Fixed Expenses = Operating income

Contribution Margin = Operating Income + Fixed Expenses

Contribution Margin = $8,000 + $32,000 = $40,000

Now, we have to calculate Variable Expenses

Sales - Variable Expenses = Contribution margin

$100,000 - Variable expenses = $40,000

Variable Expenses = $100,000 -$40,000 = $60,000

# Calculation of Units Sold.

Solution:- Contribution margin per unit $10 (given)

Contribution margin = $40,000

Units Sold = Contribution margin / Contribution margin per unit

Units sold = $40,000 / $10 = 4,000 units

Case 3. Calculation of Sales

Solution:- Sales per unit - Variable cost per unit = Contribution per unit

Sales per unit - ($70,000 / 10,000 units) = $13

Sales per unit = $13 + $7 = $20 per unit

Sales = Units * Selling price per unit

Sales = 10,000 * $20 = $200,000

# Calculation of Fixed Expenses

Solution:- Contribution Margin - Fixed Expenses = Operating income

$130,000 - Fixed Expenses = $12,000

Fixed Expenses = $130,000 -$12,000 = $118,000

Contribution Margin = Contribution margin per unit * Units sold

Contribution margin = $13 * 10,000 = $130,000

Case 4. Calculation of Contribution Margin per unit.

Solution:- Contribution Margin - Fixed Cost = Operating Loss

Contribution Margin - $100,000 = $(10,000)

Contribution Margin = $(10,000) + $100,000 = $90,000

Contribution Margin per unit = Contribution margin / Units Sold

Contribution Margin per unit = $90,000 / 6,000 = $15

# Calculation of Variable Expenses.

Solution :- Sales - Variable Expenses = Contribution Margin

= $300,000 - Variable Expenses = $90,000

Variable Expenses = $300,000 - $90,000 = $210,000

Requirement b-: Assume that more than one product is being sold in each of the following four case situation.

Case Sales Variable Expenses Contribution Margin (Percentage) Fixed Expenses Operating Income (loss)
1. $500,000

$400,000

(Working Note)

20% $93,000 (Working Note) $7,000
2. $400,000 $260,000 35% (Working Note) $100,000 $40,000 (Working Note)
3. $250,000 (Working Note) $100,000 (Working Note) 60% $130,000 $20,000
4. $600,000 $420,000 30%   (Working Note) $185,000 (Working Note) $(5,000)

Working Note :-

Case 1. Calculation of Variable Expenses.

Solution:- First of all we have to find the contribution margin.

Contribution margin (%) = Contribution margin / Sales *100%

Contribution Margin = Sales * Contribution Margin (%)

Contribution Margin = $500,000 * 20% = $100,000

Sales - Variable Expenses = Contribution Margin

$500,000 - Variable Expenses = $100,000

Variable Expenses = $500,000 - $100,000 = $400,000

# Calculation of Fixed Expenses.

Solution:- Contribution Margin - Fixed Expenses = Operating Income

$100,000 - Fixed Expense = $7,000

Fixed Expenses = $100,000- $7,000 = $93,000

Case 2. Calculation of Contribution margin (Percentage)

Solution:- Contribution margin = Sales - Variable Expenses

Contribution Margin = $400,000 - $260,000 = $140,000

Contribution Margin (%) = Contribution Margin / Sales * 100%

Contribution Margin (%) = $140,000 / $400,000 * 100% = 35%

# Calculation of Operating Income (loss).

Solution :- Contribution Margin - Fixed Expenses = Operating Income (Loss)

Operating Income = $140,000 - $100,000 = $40,000

Case 3. Calculation of Sales.

Solution :- First of all, we have to calculate Contribution Margin.

Contribution Margin - Fixed Cost = Operating Income

Contribution Margin = Operating Income + Fixed Expenses

Contribution Margin = $20,000 + $130,000 = $150,000

Contribution Margin (%) = Contribution Margin / Sales * 100 %

60% = $150,000 / Sales * 100 %

Sales = $150,000 / 60% = $250,000

# Calculation of Variable Expenses.

Solution :- Sales - Variable Expense = Contribution Margin

$250,000 - Variable Expenses = $150,000

Variable Expenses = $250,000 - $150,000 = $100,000

Case 4. Calculation of Contribution Margin (Percentage).

Solution :- First of all we have to calculate Contribution Margin.

Sales - Variable Expenses = Contribution Margin

Contribution Margin = $600,000 - $420,000 = $180,000

Contribution Margin (%) = Contribution margin / Sales * 100%

Contribution Margin (%) = $180,000 / $600,000 * 100 % = 30%

# Calculation of Fixed Expenses.

Solution :- Contribution Margin - Fixed Expenses = Operating Loss

$180,000 - Fixed Expenses = $(5,000)

Fixed Expenses = $180,000 + $ 5,000 = $185,000

Thank you..

If you need help with the another question, then you have to submit a new request for next question.

Have a Nice day Ahead.